In a hydraulic press, a force of 20N is applied to a piston of area 0.2m. if the area of the other piston is 2m^2, what is the pressure transmitted through the fluid
1
Expert's answer
2019-05-27T10:47:44-0400

The pressure of first piston:

"p_1=\frac{F_1}{S_1}"

According to Pascal's law "p_1=p_2".

Then

"\frac{F_1}{S_1}=\frac{F_2}{S_2}"

and

"F_2=\frac{F_1 S_2}{S_1}=\frac{20 \times 2}{0.2}=200N"